OpenStudy (anonymous):

please help will fan and medal ! Solve. 4x + 2 ≥ –6 A. x ≤ –2 B. x ≥ –2 C. x ≥ –1 D. x ≤ –1

OpenStudy (anonymous):

ok

OpenStudy (er.mohd.amir):

4x+2>= -6 add -2 both sides 4x+2-2 >= -6-2 4x>= -8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

will this work with any problem

OpenStudy (er.mohd.amir):

divide by 4 on both side gives x >= -8/4 x>= -2

OpenStudy (er.mohd.amir):

yes do same in all inequility

OpenStudy (anonymous):

thank u iwas very confused on the less than more than signs

OpenStudy (anonymous):

thxs

OpenStudy (er.mohd.amir):

sign is change when divide or multiply by -ve sign

OpenStudy (anonymous):

k got it

OpenStudy (er.mohd.amir):

if -x > 4 than x < -4

OpenStudy (anonymous):

k can u help with 2 more please

OpenStudy (er.mohd.amir):

yes

OpenStudy (anonymous):

Which inequality is graphed on the number line shown? A. x > –3 B. x ≥ –3 C. x < –3 D. x ≤ –3

OpenStudy (er.mohd.amir):

D x <= -3

OpenStudy (anonymous):

thannk u nver mind about the next one i figured it out thxs for all the help c u around